# Better Fair than Sorry: Adversarial Missing Sensitive Value Imputation for Fair GNNs

### v3

@@ Abstract
Graph Neural Networks (GNNs) have achieved state-of-the-art results in many relevant tasks where decisions might disproportionately impact specific communities. However, existing work on fair GNNs often assumes that either protected attributes are fully observed or that the missing protected attribute imputation is fair. In practice, biases in the imputation will propagate to the model outcomes, leading them to overestimate the fairness of their predictions. We address this challenge by proposing Better Fair than Sorry (BFtS), a fair missing data imputation model for protected attributes. The key design principle behind BFtS is that imputations should approximate the worst-case scenario for fairness—i.e. when optimizing fairness is the hardest. We implement this idea using a 3-player adversarial scheme where two adversaries collaborate against a GNN-based classifier, and the classifier minimizes the maximum bias. Experiments using synthetic and real datasets show that BFtS often achieves a better fairness $\times$ accuracy trade-off than existing alternatives.
@@ 1 Introduction
With the increasing popularity of machine learning models for high-stakes decision-making, it has become a consensus that these models carry implicit biases, and biases should be addressed to improve the fairness of algorithmic decisions Ghallab (2019). The disparate treatment of such models towards African Americans and women has been illustrated in the well-documented COMPAS Angwin et al. (2022) and Apple credit card Vigdor (2019) cases, respectively. While there has been extensive research on fair algorithms and fair machine learning, the proposed solutions have mostly disregarded important challenges that arise in real-world settings.
@@
Existing work in fair machine learning has mostly focused on tabular, image, and text data. However, in several applications, data can be naturally modeled as graphs (or networks), representing different objects, their relationships, and attributes. Graph Neural Networks (GNNs) have achieved state-of-the-art results in graph machine learning tasks, including node classification, link prediction, and graph classification Gilmer et al. (2017); Kipf and Welling (2017). For instance, in the case of link prediction in a professional network, such as LinkedIn, which is a recruiting and networking tool, link recommendations should not be biased against protected groups Fabbri et al. (2020); Current et al. (2022). However, guaranteeing fairness in graph data is a challenge due to correlations in the connections caused by homophily and influence.
@@
Another challenge for fair machine learning is the pervasiveness of missing data Little and Rubin (2019). This is particularly critical because fair algorithms often require knowledge of sensitive attributes that are more likely to be missing due to biases in the collection process or privacy concerns. As an example, the absence of data in census surveys and health surveys is correlated with sensitive attributes such as gender, age, race, etc. O’Hare and O’Hare (2019); Weber et al. (2021). Missing data is also a concern in networked data, especially when tracking the transmission of diseases Ghani et al. (1998). As a consequence, many fair machine learning algorithms require the use of missing imputation methods to mitigate their biases, and biases in imputed values might be inadvertently propagated to the final decisions Mansoor et al. (2022).
@@
This paper addresses the problem of mitigating bias in graph machine learning in the presence of missing sensitive attributes. We propose an adversarial imputation method that generates challenging instances for a broad class of fair algorithms based on GNNs. Figure 1 motivates our approach. A machine learning algorithm is applied to decide whether individuals should or should not receive credit. Figure 1(a) shows both the gender and outcome for each individual. The genders of some samples are unknown in Figure 1(a), and the demographic parity ($\Delta DP$) considering only the observed values is $0.25$ ($\Delta DP$ is a bias metric described in Section 4.2). We illustrate how different missing data imputations affect the fairness of credit decisions in Figures 1(b) and 1(c). A straightforward imputation is demonstrated in Figure 1(b), where the gender of the majority of neighbors is assigned to the missing attribute, resulting in a best-case scenario in terms of fairness with a $\Delta DP$ of $0.09$. However, in the worst-case scenario, shown in Figure 1(c), $\Delta DP$ is $0.47$. The dataset’s bias might be highly underrepresented in the imputation of Figure 1(b), and the final decisions of a fair algorithm will still be unfair in that case. We show the correlation of sensitive information and true class labels of different datasets (details in supp. material) in Figure 2. We use a Graph Convolutional Network (GCN) to predict the missing sensitive values. When $<50\%$ sensitive information is available, the predicted value underrepresented the true underlying bias in all datasets. Our method, BfTS, can better represent the true bias in the dataset. Our results provide strong evidence that off-the-shelf imputation methods often misguide fair graph machine learning algorithms due to underestimated bias.
Figure 2: Bias in training (i.e., the correlation between the sensitive attribute and node labels) for independent imputation and our approach (BFtS) with varying missing ratios. The rightmost column shows the true bias. Compared with BFtS, independent imputation often underestimates the true bias in the data.
We propose Better Fair than Sorry (BFtS), which is a 3-player adversarial learning framework for missing sensitive value imputation. Figure 3 provides an overview of the proposed approach. The GNN $f_{C}$ interacts with two adversaries $f_{A}$ and $f_{I}$. $f_{A}$ attempts to predict sensitive values based on the representations learned by $f_{C}$ while $f_{I}$ imputes missing values that help $f_{A}$ maximize its accuracy. As a high accuracy of $f_{A}$ indicates high bias in the representations learned by $f_{C}$, $f_{C}$’s parameters are optimized to minimize the bias of its representations in the worst case in terms of fairness. Our experiments show that BFtS achieves a better accuracy $\times$ fairness trade-off than existing approaches. We summarize the contributions of this paper as follows:
@@
  • We propose Better Fair than Sorry (BFtS), a novel 3-player adversarial learning framework for the imputation of missing sensitive data that produces worst-case imputed values for fair GNNs and is effective even when sensitive attribute information is completely unavailable.
  • We prove that the maximal demographic parity (Eq. 8) of BFtS is minimal among all estimators of missing sensitive attributes and that BFtS is robust to convergence issues in adversarial training.
  • We show empirically that BFtS achieves a better fairness $\times$ accuracy trade-off than the existing alternatives.
@@ 1.1 Related Work
Fairness in graph machine learning. We divide the work in this area into three main directions (pre-processing, extended objective functions, and adversarial learning). FairOT, FairDrop, FairSIN, EDITS, and Graphair are pre-processing techniques that remove the bias from the graph before training Laclau et al. (2021); Spinelli et al. (2021); Yang et al. (2024); Dong et al. (2022); Ling et al. (2023). Fairwalk, Crosswalk, Debayes, MONET, NIFTY, FairVGNN, PFR-AX, and FairGAE modify the objective function of GNNs to learn fair representations during training Rahman et al. (2019); Khajehnejad et al. (2022); Buyl and De Bie (2020); Palowitch and Perozzi (2019); Agarwal et al. (2021); Wang et al. (2022); Merchant and Castillo (2023); Fan et al. (2021). CFC, FLIP, DKGE, and Debias train adversarial networks along with GNN for fair prediction Bose and Hamilton (2019); Masrour et al. (2020); Arduini et al. (2020); Zhang et al. (2018). However, different from our work, all these methods require fully observed sensitive attributes to ensure fairness.
@@
Missing Data Imputation and fairness. Missing data can be imputed using unconditional mean, reconstruction, preferential attachment, autoencoders, etc. Donders et al. (2006); Pereira et al. (2020); Huisman (2009). The traditional procedure for handling missing data is to first impute the missing values, and then solve the task Rossi et al. (2021); Buck (1960). SAT is a matching-based GNN for graphs with missing attributes Chen et al. (2020), but it does not account for fairness. Training a classifier from imputed data can amplify the unfairness of a machine learning model, as discussed in Subramonian et al. (2022); Zhang and Long (2021); Feng et al. (2023). Some studies try to generate fair feature imputations. Subramonian et al. (2022) has modified mean aggregation feature imputation to minimize discrimination risk against protected groups. Feng et al. (2023) has proposed recursive partitioning and fair uniform resampling for fair classification with missing values. However, as the approaches discussed earlier, these methods require full knowledge of the protected attributes.
@@
Fairness with missing protected attributes. These studies are the ones most related to ours. One extreme assumption is the complete unavailability of sensitive information. Hashimoto et al. (2018) takes into account the worst-case distribution across a given group size, and Lahoti et al. (2020) adversarially reweighs training samples to optimize the average loss. Chai and Wang (2022) employs top $k$ average loss minimization to guarantee fairness. Zhao et al. (2022) minimizes the correlation between the predictions and other features associated with sensitive attributes. Yan et al. (2020) achieves fairness via class re-balancing and Chai et al. (2022) uses soft labels from an overfitted teacher model. In practice, some sensitive information is often available and can significantly improve fairness Chai and Wang (2022). FairGNN Dai and Wang (2021) assumes limited sensitive information and learns a fair GNN by predicting the missing values with an independent imputation model. RNF Du et al. (2021) generates proxy annotations for missing attributes using generalized cross-entropy Zhang and Sabuncu (2018). Similar to RNF, our approach (BFtS) handles completely or partially missing sensitive information. BFtS applies a 3-player scheme to minimize the maximum possible bias and outperforms FairGNN and RNF in terms of fairness $\times$ accuracy. Minimizing the maximum instead of the average risk has been shown to achieve better guarantees Shalev-Shwartz and Wexler (2016). A similar minmax approach has been applied to maximize the robustness and accuracy of uncertainty models Löfberg (2003); Lanckriet et al. (2002); Chen et al. (2017); Fauß et al. (2021). In Nguyen et al. (2017); Vandenhende et al. (2019), a 3-player adversarial network is proposed to improve the classification and stability of adversarial learning.
2 Problem Definition and Preliminaries
@@
In this section, we formalize the problem addressed in this paper and introduce the Graph Neural Network model, which is our base model for node classification. For simplicity, we only consider binary-sensitive attributes. By modifying the loss function in Eq. 2 to categorical cross entropy and MSE, BFtS can be generalized for non-binary and continuous protected attributes, respectively.
@@ 2.1 Notation and Problem Definition
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X},\mathcal{S})$ be an undirected graph where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, $\mathcal{X}$ are node attributes, and $\mathcal{S}$ is the sensitive attribute. The matrix $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of $\mathcal{G}$ where $A_{uv}=1$ if there is an edge between $u$ and $v$ and $A_{uv}=0$, otherwise. We focus on the setting where sensitive attributes might be missing and only $\mathcal{V}_{S}\subseteq\mathcal{V}$ nodes include the information of the sensitive attribute $s_{v}$ for a node $v$. The sensitive attribute forms two groups, which are often called sensitive ($s_{v}=1$) and non-sensitive ($s_{v}=0$).
@@
$\min_{f_{C}}\mathcal{L}_{C}(f_{C})+\alpha\mathcal{L}_{A}(f_{C})$
@@
Computing fairness loss $\mathcal{L}_{A}$ requires sensitive attributes $s_{v}$. The main challenge addressed in this paper is due to our assumption that sensitive attributes can be missing. In this scenario, one can apply an independent imputation model $s_{v}\approx f_{I}(X[v])$ to predict missing sensitive attributes before training the fair classifier. However, the fairness of the resulting classifier will be highly dependent on the accuracy of $f_{I}$, which is the key motivation for our work. In particular, we consider two scenarios to motivate our solution:
Sensitive attributes missing completely at random (MCR) at a high rate:
The missingness process can be modeled by a Bernoulli distribution $Bern(\beta)$ with $\beta$ close to $1$ and thus $f_{I}$ is expected to have a high error rate. Thus, the bias $\mathcal{L}_{A}$ can be underestimated by the fair classification model.
Adversarily missing sensitive attributes:
The missingness process is generated by an adversary that manipulates $f_{I}$ to artificially decrease the fairness loss $\mathcal{L}_{A}$. As an example, the adversary might select missing values to flip the predictions of $s_{v}$ for nodes that are both in the protected group and the undesired class ($y=0$). These scenarios are particularly relevant in graphs as they can exploit structural graph properties (see Sec. 4.1).
As an independent imputation cannot be trusted in these scenarios, we propose an interactive (3-player) worst-case imputation model for fair classification described in Sec. 3.
2.2 Graph Neural Network
Two of the models (or players) learned by our model are GNNs. Without loss of generality, we will assume that they are message-passing neural networks Gilmer et al. (2017), which we will simply refer to as GNNs. A GNN learns node representations $\mathbf{h}_{v}$ for each node $v$ that can be used to predict (categorical or continuous) values $y_{v}=f(\mathbf{h}_{v})$. These representations are learned via successive $\mathsf{AGGREGATE}$ and $\mathsf{COMBINE}$ operations defined as follows:
Figure 3: Proposed 3-player framework for fair GNN training with missing data imputation (BFtS). $f_{C}$ generates node representations by minimizing the classification loss $\mathcal{L}_{C}$ (Eqn. 1) and maximizing sensitive attribute prediction loss $\mathcal{L}_{A}$ (Eqn. 3). $f_{A}$ predicts sensitive attributes using the representations from $f_{C}$ by minimizing $\mathcal{L}_{A}$. $f_{I}$ predicts missing values by minimizing the imputation loss $\mathcal{L}_{I}$ (Eqn. 2) and maximizing $\mathcal{L}_{A}$. $\hat{y}$, $\hat{si}$, $\hat{sa}$ are the predictions generated by $f_{C}$, $f_{I}$ and $f_{A}$, respectively.

### v4

# Fair Graph Machine Learning under Adversarial Missingness Processes

@@ Abstract
Graph Neural Networks (GNNs) have achieved state-of-the-art results in many relevant tasks where decisions might disproportionately impact specific communities. However, existing work on fair GNNs often assumes that either sensitive attributes are fully observed or they are missing completely at random. We show that an adversarial missingness process can inadvertently disguise a fair model through the imputation, leading the model to overestimate the fairness of its predictions. We address this challenge by proposing Better Fair than Sorry (BFtS), a fair missing data imputation model for sensitive attributes. The key principle behind BFtS is that imputations should approximate the worst-case scenario for fairness—i.e. when optimizing fairness is the hardest. We implement this idea using a 3-player adversarial scheme where two adversaries collaborate against a GNN classifier, and the classifier minimizes the maximum bias. Experiments using synthetic and real datasets show that BFtS often achieves a better fairness $\times$ accuracy trade-off than existing alternatives under an adversarial missingness process.
@@ 1 Introduction
With the increasing popularity of machine learning in high-stakes decision-making, it has become a consensus that these models carry implicit biases, and these biases should be addressed to improve the fairness of algorithmic decisions ghallab2019responsible . The disparate treatment of such models towards African Americans and women has been illustrated in the well-documented COMPAS b1 and Apple credit card apple cases, respectively. While there has been extensive research on fair machine learning, the proposed solutions have mostly disregarded important challenges that arise in real-world settings. For instance, in many applications, data can be naturally modeled as graphs (or networks), representing different objects, their relationships, and attributes instead of as sequences or images.
@@
Another challenge for fair machine learning is the pervasiveness of missing data little2019statistical . This is particularly critical because fair algorithms often require knowledge of sensitive attributes that are more likely to be missing due to biases in the collection process or privacy concerns. As an example, the absence of data in census surveys and health surveys is correlated with sensitive attributes such as gender, age, race, etc. o2019potential ; weber2021gender . Missing data is also a concern in networked data, especially when tracking the transmission of diseases ghani1998sampling . As a result, many fair machine learning algorithms rely on imputation methods to handle missing data and mitigate associated biases. However, when the missingness mechanism is not random, standard imputation techniques can introduce systematic errors that propagate through the pipeline and compromise downstream fairness mansoor2022impact ; jeong2022fairness .
This work investigates the impact of adversarial sensitive value missingness processes on fairness, where the pattern of missing sensitive data is structured to obscure true disparities. If the predictions from the imputation method fail to capture the true distribution of a sensitive attribute, any fairness-aware model trained on the imputed data is prone to inheriting the underlying true bias. More specifically, a key challenge arises when an adversarial missingness process makes the imputed dataset appear to be fair. In graphs, an adversary can exploit the graph structure to manipulate the imputed values and, as a consequence, the fairness-aware model. Prior work on the fairness of graph machine learning assumes sensitive values are Missing Completely At Random (MCAR). However, this assumption rarely holds in practice o2019potential ; weber2021gender ; ghani1998sampling ; jeong2022fairness . Methods that overlook the adversarial missingness process risk producing misleading fairness guarantees w.r.t. the complete data.
@@
Figure 1 motivates our problem setting using a toy example. A machine learning algorithm is applied to decide whether individuals should or should not receive credit. Figure 1(a) shows both the gender and outcome for each individual. The genders of some samples are unknown in Figure 1(a), and the demographic parity ($\Delta DP$) considering only the observed values is $0.25$ ($\Delta DP$ is a bias metric described in Section 5). We illustrate how different missing data imputations affect the fairness of credit decisions in Figures 1(b) and 1(c). A straightforward imputation is demonstrated in Figure 1(b), where the gender of the majority of neighbors is assigned to the missing attribute, resulting in a best-case scenario in terms of fairness with a $\Delta DP$ of $0.09$. However, in the worst-case scenario, shown in Figure 1(c), $\Delta DP$ is $0.47$. If an adversarial missingness process can induce the imputation model to generate the imputation from Figure 1(b) but the complete data is Figure 1(c), a fair model trained on the imputed data will still be biased w.r.t. the complete data.
To counter an adversarial missingness process for sensitive values on graphs, we propose Better Fair than Sorry (BFtS), a 3-player adversarial learning framework for missing sensitive value imputation based on Graph Neural Networks (GNNs). Our experiments show that BFtS achieves a better accuracy $\times$ fairness trade-off than existing approaches, especially under adversarial missing sensitive values.
@@
We summarize the contributions of this paper as follows: (1) We investigate theoretically and empirically the potential of an adversarial missingness process to bias a fair graph machine learning model; (2) We propose Better Fair than Sorry (BFtS), a novel 3-player adversarial learning framework for the imputation of missing sensitive data that produces worst-case imputed values for fair GNNs and is effective under adversarial missingness processes and even when sensitive attribute information is completely unavailable; and (3) We show empirically that BFtS achieves a better fairness $\times$ accuracy trade-off than the existing alternatives.
@@ 1.1 Related Work
Fairness in graph machine learning. We divide the work in this area into three main directions (pre-processing, extended objective functions, and adversarial learning). FairOT, FairDrop, FairSIN, EDITS, and Graphair are pre-processing techniques that remove the bias from the graph before training fairot ; fairdrop ; fairsin ; edits ; graphair . Fairwalk, Crosswalk, Debayes, MONET, NIFTY, FairVGNN, PFR-AX, FairSAD, and FairGAE modify the objective function of GNNs to learn fair representations during training fairwalk ; crosswalk ; debayes ; monet ; agarwal2021towards ; wang2022improving_short ; merchant2023disparity ; zhu2024fair ; fan2021fair . CFC, FLIP, DKGE, and Debias train adversarial networks along with GNN for fair prediction cfc ; cfc2 ; kg ; zhang2018mitigating . However, different from our work, all these methods require fully observed sensitive attributes to incorporate fairness considerations.
@@
Missing Data Imputation and fairness. Missing data can be imputed using unconditional mean, reconstruction, preferential attachment, autoencoders, etc. donders2006gentle ; pereira2020reviewing ; huisman2009imputation . The traditional procedure for handling missing data is independent imputation—i.e., to first impute the missing values and then solve the task rossi2021unreasonable ; buck1960method . SAT is a matching-based GNN for graphs with missing attributes chen2020learning , but it does not account for fairness. Training a classifier from imputed data can amplify the bias of a machine learning model, as discussed in subramoniandiscrimination ; zhang2021assessing ; feng2023adapting ; guha2024automated ; martinez2019fairness ; fricke2020missing ; mansoor2022impact . Some studies try to generate fair feature imputations subramoniandiscrimination ; feng2023adapting ; jeong2022fairness ; zhang2022fairness . However, as the approaches discussed earlier, these methods do not consider missing sensitive information and require full knowledge of the sensitive attributes for fairness intervention.
@@
Fairness with missing sensitive attributes. These studies are the most related to ours. One extreme assumption is the complete unavailability of sensitive information. For instance, hashimoto2018fairness takes into account the worst-case distribution across a given group size, and lahoti2020fairness adversarially reweighs training samples to optimize the average loss. In chai2022self , a top $k$ average loss minimization scheme is applied for fairness. In zhao2022towards , the correlation between the predictions and other features associated with sensitive attributes is minimized. Other approaches to achieve fairness are class re-balancing yan2020fair and chai2022fairness using soft labels from an overfitted teacher model. However, in practice, some sensitive information is often available and can significantly improve fairness chai2022self . The above-mentioned models fail to utilize the accessible sensitive values. FairGNN fairgnn assumes limited sensitive information and learns a fair GNN by predicting the missing values with an independent imputation model. FairAC guo2023fair assumes that both node features and the sensitive attribute are missing for a subset of nodes, and utilizes only the observed sensitive attributes without applying any imputation. RNF du2021fairness generates proxy annotations for missing attributes using generalized cross-entropy zhang2018generalized . Similar to RNF, our approach (BFtS) handles completely or partially missing sensitive information. BFtS applies a 3-player scheme to minimize the maximum possible bias and outperforms FairGNN and RNF in terms of fairness $\times$ accuracy. Minimizing the maximum instead of the average risk, similar to our method, has been shown to achieve better guarantees shalev2016minimizing . A similar minmax approach has been applied to maximize the robustness and accuracy of uncertainty models lofberg2003minimax ; lanckriet2002robust ; chen2017robust ; fauss2021minimax . In nguyen2017dual ; vandenhende2019three , a 3-player adversarial network is proposed to improve the classification and stability of adversarial learning.
Adversarial attacks on fairness. Recent work has investigated data poisoning as a means to adversarially degrade model fairness mehrabi2021exacerbating ; solans2020poisoning . Some approaches, such as UnfairTrojan and TrojFair introduce backdoor attacks specifically aimed at reducing model fairness furth2024unfair ; zheng2023trojfair . In the context of graph neural networks (GNNs), adversarial fairness attacks have predominantly focused on modifying graph topology. For instance, hussain2022adversarial ; zhang2023adversarial alter the structure of the graph to compromise the fairness of GNN predictions. NIFA luo2024your attacks GNN fairness by injecting new nodes guided by an uncertainty-maximization and homophily-enhancement strategy. Similarly, FATE kang2023deceptive employs a bilevel optimization-based meta-learning framework to carry out poisoning attacks on GNNs. These existing approaches primarily target fairness through data poisoning, node injection, or structural perturbations. To the best of our knowledge, no prior work has explored attacks that can only manipulate the missingness process for sensitive attributes as a means to compromise model fairness.
2 Preliminaries
@@
In this section, we formalize the setting considered in this paper and introduce the Graph Neural Network model, which is our base model for node classification. For simplicity, we only consider binary-sensitive attributes. By modifying the loss function in Eq. 2 to categorical cross entropy and MSE, our work can be generalized for non-binary and continuous sensitive attributes, respectively.
@@ 2.1 Notation and Problem Definition
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X},\mathcal{S})$ be an undirected graph where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, $\mathcal{X}$ are node attributes, and $\mathcal{S}$ is the set of sensitive attributes. The matrix $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of $\mathcal{G}$ where $A_{uv}=1$ if there is an edge between $u$ and $v$ and $A_{uv}=0$, otherwise. We focus on the setting where sensitive attributes might be missing and only $\mathcal{V}_{S}\subseteq\mathcal{V}$ nodes include the information of the sensitive attribute $s_{v}$ for a node $v$. The sensitive attribute forms two groups, which are often called sensitive ($s_{v}=1$) and non-sensitive ($s_{v}=0$).
@@
$f_{C}^{*}=\arg\min_{f_{C}}\mathcal{L}_{C}(f_{C})+\alpha\mathcal{L}_{A}(f_{C})$
@@
Computing fairness loss $\mathcal{L}_{A}$ requires sensitive attributes $s_{v}$. The main challenge addressed in this paper is due to our assumption that sensitive attributes can be missing. In this scenario, one can apply an independent imputation model $s_{v}\approx f_{I}(v)$ to predict missing sensitive attributes before training the fair classifier. However, the fairness of the resulting classifier will be highly dependent on the accuracy of $f_{I}$, which is the key motivation for our work. Two of the models (or players) learned by our model are GNNs. The details about Graph Neural Networks are provided in Appendix A.
3 Introducing Bias via an Adversarial Missingness Process
Our work investigates the role of missing sensitive values in the fairness of graph machine learning. Datasets with prevalent missing values often require the application of a missing data imputation method before they can be processed by a fair machine learning model. In this section, we focus on how the missingness process (i.e., the process that generates missing values) can lead to (intended or unintended) biases in a fair model trained using the imputed data. More specifically, the missingness process can induce the fair model to underestimate the bias in the imputed data. As a consequence, the fair model applied to the imputed data can still be inadvertently biased w.r.t. the complete data.
The simplest missingness process for sensitive values is missing completely at random—i.e., the probability of a value being missing is independent of the data. However, in practical scenarios, missing values are rarely random o2019potential ; weber2021gender ; ghani1998sampling . We focus on the particular case where the missingness process is adversarial and unknown. We prove that manipulating the missingness process to induce bias optimally is computationally hard. However, we show empirically that a simple heuristic can effectively introduce bias to an independent imputation model that simply tries to maximize imputation accuracy. We formulate the problem for a given number of observed sensitive values $k$.
Definition 1.
Adversarial Missingness for Minimizing Classification Bias (AMMCB): Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with a feature matrix $\mathcal{X}$, hypothesis spaces $\Theta_{C}$ and $\Theta_{I}$ and a training dataset $\mathcal{V}_{L}$, select a set of nodes $\mathcal{V}_{S}^{*}\subset\mathcal{V}$ such that $|\mathcal{V}_{S}^{*}|=k$ as to minimize the fairness loss $\mathcal{L}_{A}$ of the optimal classifier $f_{C}^{*}\in\Theta_{C}$ when it is trained with sensitive values imputed from the optimal imputation model $f_{I}^{*}$ that itself is trained with $\mathcal{V}_{S}^{*}$.
AMMCB is a discrete tri-level optimization problem as $f_{I}^{*}$ depends on $\mathcal{V}_{S}^{*}$ and $f_{C}^{*}$ depends on $f_{I}^{*}$. We note that by minimizing the bias loss $\mathcal{L}_{A}$, the adversary prevents the fair classifier from reducing the bias in its predictions w.r.t. the complete data. Moreover, AMMCB requires knowledge of the hypothesis spaces $\Theta_{C}$ and $\Theta_{I}$, which might not be accessible by the adversary in practice. It is still convenient to formalize AMMCB because it defines the ultimate goal of the adversary. However, instead of solving this problem, we focus on the following more tractable surrogate.
Definition 2.
Adversarial Missingness for Minimizing the Bias in the Data (AMMBD): Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with a feature matrix $\mathcal{X}$, a hypothesis space $\Theta_{I}$, and a training dataset $\mathcal{V}_{L}$, select $\mathcal{V}_{S}\subset\mathcal{V}$ such that $|\mathcal{V}_{S}|=k$ as to minimize the dataset bias generated by an imputation model $f_{I}^{*}\in\Theta$ over $\mathcal{V}_{L}$.
AMMBD does not depend on $\Theta_{C}$ and its objective is constrained to the training data. The relationship between AMMBD and AMMCB is due to generalization—i.e., the expectation that the bias in $\mathcal{V}_{L}$ will generalize to the predictions from $f_{C}$. However, we show that even AMMDB is NP-hard.
Theorem 1.
The AMMBD problem is NP-hard.
See proof of Theorem 1 in Appendix A. The idea of our proof is to frame AMMBD as an adversarial version of active learning where the goal is to strategically minimize the accuracy of the model for certain nodes by selecting labels. More specifically, we apply a popular formulation of active learning as a coverage problem yehuda2022active ; ren2021survey . The adversary will target nodes either (1) with sensitive attribute $s_{v}=1$ and in the undesired class $y_{v}=0$ or (2) with sensitive attribute $s_{v}=1$ and in the desired class $y_{v}=1$. Targeting these nodes will decrease the bias in the imputed data.
Theorem 1 can be interpreted as a positive result, as it shows that, in theory, a simpler surrogate of the adversary’s objective is hard to optimize. The best-known approximation for the minimum k-union problem chlamtavc2017minimizing used in our reduction is $O(n^{1/4+\epsilon})$, where $n$ is the number of sets, which is mapped to the size of the graph in our problem. This means that approximately $k/n^{1/4}$ nodes would have to be selected to achieve the optimal coverage for $k$ nodes. However, we show empirically that a simple heuristic based on node degrees often achieves good results for the problem when an independent imputation model $f_{I}$ based on a GNN is applied.
Heuristic for AMMBD: We generate missing values synthetically in our experiments. To simulate an adversarial missingness process, we select low-degree nodes to have missing values. This choice is based on previous work showing that low-degree nodes have lower prediction accuracy tang2020investigating ; liu2023generalized ; ju2024graphpatcher . To generate $k$ nodes with missing sensitive data, we choose the $k$ nodes with the lowest degree.
To see the impact of adversarial missingness, we compare the imputation of adversarial missingness and random missingness. We show the correlation of sensitive information and true class labels for different datasets in Figure 2. We use a Graph Convolutional Network (GCN) to predict the missing sensitive values for independent imputation. Figures 2(a) and 2(b) show the correlation with adversarial missingness and random missingness, respectively. When $50\%$ sensitive information is available, the predicted value with independent imputation underrepresents the true underlying bias in all datasets with adversarial missingness. For random missingness, the performance is slightly better. Figure 2 (c) and (d) show results for BFtS, which will be introduced in the next section. BFtS rarely underestimates the true bias in the datasets. Our results provide strong evidence that off-the-shelf imputation methods often misguide fair graph machine learning algorithms due to underestimated bias, especially with adversarial missingness processes.
Figure 2: Empirical results showing that a simple degree-based adversarial missingness process is effective at minimizing the bias of an independent imputation model (GNN) compared with a random missingness process. The last column of each matrix shows the true bias in the data, and lower values show that the bias is underestimated. We also show results for our approach (BFtS) described in Section 4 that addresses this problem using a 3-player adversarial imputation method.
As our proposed model operates under a worst-case bias assumption, it may overestimate the bias in the complete data, as illustrated in Figure 2 using the Bail dataset. This results in a trade-off between fairness and utility, which is governed by a hyperparameter $\beta$. Figures 3(a), (b), and (c) show the F1 score, $1-\Delta\text{DP}$, and $1-\Delta\text{EQOP}$, respectively, as $\beta$ varies and with $30\%$ of sensitive values observed. Here, F1 serves as the utility metric, while $1-\Delta\text{DP}$ and $1-\Delta\text{EQOP}$ quantify fairness (see Section 5). We compare the performance of our BFtS model against a fair adversarial model trained with complete data. Adjusting $\beta$ enables controlling the balance between utility and robustness to worst-case bias, with appropriate tuning yielding an effective trade-off.
Figure 3: Fairness–utility trade-off of BFtS on the Bail dataset as $\beta$ varies, compared to a fair adversarial model trained with complete data. Lower $\beta$ yields higher accuracy but lower fairness.